### v1

# CP Violation in $D\to KK$ Decays: A Comparative Analysis of Triplet and Sextet Diquarks

Abstract
Recent measurements of the CP asymmetry in the decay $D^{0}\rightarrow K_{S}^{0}K_{S}^{0}$ by the CMS collaboration, $A_{CP}(K_{S}^{0}K_{S}^{0})=(6.2\pm 3.0\pm 0.2\pm 0.8)\%$, and by LHCb, $A_{CP}(D^{0}\to K_{S}^{0}K_{S}^{0})=(1.86\pm 0.23\pm 0.11)\%$, suggest possible deviations from Standard Model (SM) expectations, which predict asymmetries below the percent level. This singly Cabibbo-suppressed decay is particularly sensitive to new physics, as the leading amplitudes vanish in the exact U-spin symmetry limit and the process is dominated by W-exchange topologies. We investigate scalar diquark contributions to this decay, comparing color-sextet and color-triplet representations. We find that the color-sextet diquark, characterized by a symmetric color structure $(C_{1}^{\mathrm{NP}}=C_{2}^{\mathrm{NP}})$, avoids color suppression and can generate CP asymmetries in the range $0.5\%$$1.5\%$ for a diquark mass of order 1 TeV. In contrast, the color-triplet contribution is strongly suppressed due to destructive interference from its antisymmetric color structure. We further show that a flavor hierarchy in the sextet couplings, with $\lambda_{ud}>\lambda_{us}$, can simultaneously account for the observed deviation from the U-spin sum rule in $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$ and the measured CP asymmetry in $D^{0}\to K_{S}^{0}K_{S}^{0}$. These results identify color-sextet scalar diquarks as viable candidates for explaining enhanced CP violation in charm decays.
I Introduction
Charge–parity (CP) violation is one of the most profound phenomena in particle physics, playing a central role in our understanding of the matter–antimatter asymmetry of the Universe. Within the Standard Model (SM), CP violation originates from a single irreducible complex phase in the Cabibbo–Kobayashi–Maskawa (CKM) quark mixing matrix. While this mechanism successfully explains the observed CP violation in the kaon and beauty sectors, its effects in the charm sector are predicted to be highly suppressed. This suppression arises from the hierarchical structure of the CKM matrix, the Glashow–Iliopoulos–Maiani (GIM) mechanism, and the dominance of light quark contributions in charm transitions. As a consequence, direct CP asymmetries in singly Cabibbo-suppressed (SCS) charm decays are generally expected to be at most of order $10^{-3}$ within the SM.
The experimental study of CP violation in charm decays has recently entered a new era of precision, driven by the large datasets collected by the LHCb and CMS experiments. Of particular interest is the decay channel $D^{0}\to K_{S}^{0}K_{S}^{0}$, which provides a uniquely sensitive probe of new sources of CP violation. Recent measurements by the LHCb collaboration have reported a direct CP asymmetry [1]:
$A_{CP}(D^{0}\to K_{S}^{0}K_{S}^{0})=(1.86\pm 0.23\pm 0.11)\%,$
while CMS has reported a consistent result at the percent level. These values are significantly larger than the SM expectation, which is usually estimated to lie below $0.5\%$ due to the strong suppression of CP-violating amplitudes in this channel[2].
The decay $D^{0}\to K_{S}^{0}K_{S}^{0}$ is a very sensitive probe for New Physics (NP)[3]. In the SM, the leading tree-level amplitude vanishes in the exact U-spin limit, leaving the decay to proceed through suppressed U-spin-breaking effects and loop-induced penguin diagrams[4, 5]. This inherent suppression minimizes the “background” of SM CP violation, making the channel an ideal laboratory for identifying new sources of CP violation[3].
Theoretical interest in charm CP violation has been further stimulated by recent experimental indications of deviations from U-spin sum rules relating the CP asymmetries of the SCS decay modes $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$. Global fits to experimental data suggest a violation of these relations at the level of approximately $2.7\sigma$, with both asymmetries observed to be positive, contrary to the leading-order SM expectation [6]. These observations point toward the intriguing possibility of new physics contributions to charm decay amplitudes.
Among the various BSM scenarios proposed to explain enhanced CP violation in charm decays, scalar diquarks provide a particularly compelling framework . Scalar diquarks are easily introduced in many well-motivated extensions of the SM, including grand unified theories [7]. based on gauge groups such as $SO(10)$ and $E_{6}$, composite models, and theories with extended color sectors. These states transform as colored scalars under $SU(3)_{C}$ and can couple directly to quark pairs, generating new four-quark interactions at tree level. Unlike SM contributions, which generate primarily vector and axial-vector operators, scalar diquark exchange can produce scalar and tensor operators, introducing new possible interference terms and CP-violating phases.
A key feature of scalar diquarks is how they transform under the color gauge group, which determines their interference properties with SM amplitudes. Specifically, scalar diquarks transforming as color sextets possess a symmetric color structure that allows constructive interference with SM exchange amplitudes, while color-triplet diquarks, characterized by antisymmetric color contractions, generating destructive interferences and are therefore less effective in enhancing CP asymmetries. This difference between sextet and triplet scalar diquarks has important phenomenological consequences for charm decays dominated by exchange topologies, such as $D^{0}\to K_{S}^{0}K_{S}^{0}$.
In this work, we perform a detailed and systematic investigation of scalar diquark contributions to CP violation in charm decays, with a primary focus on the $D^{0}\to K_{S}^{0}K_{S}^{0}$ channel. We consider both color-sextet and color-triplet scalar diquark representations and their effects on the effective weak Hamiltonian governing charm transitions is studied. We derive the resulting Wilson coefficients, examine the induced operator structures, and estimate the resulting CP asymmetries. The implications of flavor-dependent diquark couplings and their ability to explain the observed pattern of CP violation across multiple charm decay channels is explored.
Our analysis demonstrates that color-sextet scalar diquarks with masses in the TeV range and perturbative couplings is able to generate CP asymmetries at the percent level, within the range of current experimental measures. In contrast, the color-triplet representation is suppressed due to destructive color interference. We also show that a hierarchical structure of diquark couplings can account for the observed violation of U-spin sum rules and provide an explanation of CP violation in both neutral and charged kaon final states.
This paper is organized as follows. In Section II, we review the Standard Model contributions to the decay $D^{0}\to K_{S}^{0}K_{S}^{0}$ and discuss the origin of CP violation in this channel. In Section III, we introduce the scalar diquark framework and derive the effective operators generated by diquark exchange. Section IV presents a detailed phenomenological analysis of scalar diquark contributions to CP asymmetries in charm decays, including a comparison between color-sextet and color-triplet scenarios. Finally, Section V summarizes our results and discusses their implications for future experimental searches and theoretical developments.
II Standard Model contribution
In the Standard Model (SM), the decay $D^{0}\rightarrow K_{S}^{0}K_{S}^{0}$ is a Singly Cabibbo-Suppressed (SCS) process. Unlike Cabibbo-favored decays which are dominated by tree-level spectator diagrams, this channel proceeds primarily through non-spectator topologies: the $W$-exchange ($E$) and Penguin Annihilation ($PA$) diagrams[3]. The effective Hamiltonian for the SCS charm decay is[4, 8]:
$\mathcal{H}_{eff}=\frac{G_{F}}{\sqrt{2}}\sum_{q=d,s}\lambda_{q}\left(C_{1}Q_{1}^{q}+C_{2}Q_{2}^{q}+\sum_{i=3}^{6}C_{i}Q_{i}\right)$
[…]
$Q_{1}^{q}$ $=$ $(\bar{u}_{\alpha}\gamma^{\mu}P_{L}q_{\beta})(\bar{q}_{\beta}\gamma_{\mu}P_{L}c_{\alpha})$ (1)
$Q_{2}^{q}$ $=$ $(\bar{u}_{\alpha}\gamma^{\mu}P_{L}q_{\alpha})(\bar{q}_{\beta}\gamma_{\mu}P_{L}c_{\beta})$ (2)
$Q_{3}$ $=$ $(\bar{u}_{\alpha}\gamma^{\mu}P_{L}c_{\alpha})\sum_{q^{\prime}}(\bar{q}^{\prime}_{\beta}\gamma_{\mu}P_{L}q^{\prime}_{\beta})$ (3)
$Q_{4}$ $=$ $(\bar{u}_{\alpha}\gamma^{\mu}P_{L}c_{\beta})\sum_{q^{\prime}}(\bar{q}^{\prime}_{\beta}\gamma_{\mu}P_{L}q^{\prime}_{\alpha})$ (4)
$Q_{5}$ $=$ $(\bar{u}_{\alpha}\gamma^{\mu}P_{L}c_{\alpha})\sum_{q^{\prime}}(\bar{q}^{\prime}_{\beta}\gamma_{\mu}P_{R}q^{\prime}_{\beta})$ (5)
$Q_{6}$ $=$ $(\bar{u}_{\alpha}\gamma^{\mu}P_{L}c_{\beta})\sum_{q^{\prime}}(\bar{q}^{\prime}_{\beta}\gamma_{\mu}P_{R}q^{\prime}_{\alpha})$ (6)
The usual $O_{7-10}$ are generated in SM through electroweak loop penguin and are strongly suppressed. It is why we should not include them in that section. Using standard notation, the full amplitude $A(D^{0}\to K_{S}^{0}K_{S}^{0})$ is expressed as[3]:
$A(D^{0}\to K_{S}^{0}K_{S}^{0})=\frac{G_{F}}{\sqrt{2}}\left[\lambda_{sd}A_{sd}-\frac{\lambda_{b}}{2}A_{b}\right]$ (7)
with $\lambda_{sd}\equiv(\lambda_{s}-\lambda_{d})/2$
$A_{sd}=\sum_{i=1,2}C_{i}\left(\langle K_{S}K_{S}|Q_{i}^{s}|D^{0}\rangle-\langle K_{S}K_{S}|Q_{i}^{d}|D^{0}\rangle\right)$ (8)
The CP-Violating Piece ($A_{b}$) is given by:
$A_{b}=\sum_{i=1}^{2}C_{i}\langle Q_{i}^{u}\rangle+\sum_{i=3}^{6}C_{i}\langle Q_{i}\rangle$ (9)
  • Term 1 ($\sum_{i=1}^{2}C_{i}\langle Q_{i}^{u}\rangle$): Represents the Tree/Exchange diagrams. It represents the ”flavor-symmetric” part of the exchange [^1] While the physical decay involves strange ($s$) and down ($d$) quarks, theorists use the $u$-quark operator as a reference point to isolate the flavor-symmetric part of the exchange diagram[3].
  • Term 2 ($\sum_{i=3}^{6}C_{i}\langle Q_{i}\rangle$): Represents the QCD Penguin loops.
Table 1: Effective Hamiltonian Operator Groups within the Standard Model for $D^{0}\to K_{S}^{0}K_{S}^{0}$. The operator basis and their phenomenological roles follow the standard $\Delta C=1$ framework [10, 4, 8]
Operator Group Type Physical Process Role in $D^{0}\to K_{S}^{0}K_{S}^{0}$
$Q_{1},Q_{2}$ Current-Current $W$-Exchange Dominant; determines the Branching Ratio.
$Q_{3}\dots Q_{6}$ QCD Penguins Gluon Loops Main source of SM CP Violation.
$Q_{7}\dots Q_{10}$ EW Penguins Photon/$Z$ Loops Isospin-violating; sensitive to heavy NP.
Neglecting higher-order terms in the ratio $|\lambda_{b}/\lambda_{sd}|$, the direct CP asymmetry is computed as [3]:
$a_{CP}^{dir}=\text{Im}\left(\frac{\lambda_{b}}{\lambda_{sd}}\right)\cdot\text{Im}\left(\frac{A_{b}}{A_{sd}}\right)$
  • Weak Factor: The first term, $\text{Im}\left(\frac{\lambda_{b}}{\lambda_{sd}}\right)\approx-6\times 10^{-4}$, is a fixed Standard Model value determined by CKM matrix elements.
  • Strong Factor: The second term, $\text{Im}\left(\frac{A_{b}}{A_{sd}}\right)=\left|\frac{A_{b}}{A_{sd}}\right|\sin\delta$, depends on the magnitude of the amplitude ratio and the relative strong phase ($\delta$) between them. In reference [3], they suggested a ratio of $|A_{b}/A_{sd}|\approx 3\text{--}5$.
In SM, a numerical estimate of the direct CP violating asymmetry in this channel is given by
$|a_{CP}^{dir}|\approx 6\times 10^{-4}\times(3\text{ to }5)\times\sin\delta\approx\mathbf{0.2\%\text{ to }0.4\%}$
III Scalar Diquarks in Beyond the Standard Model Theories
In this work, we investigate the impact of scalar diquark exchange as a concrete and well-motivated NP scenario capable of generating sizeable CP-violating effects in this channel.
The classification of scalar diquarks based on their transformation properties under the Standard Model gauge group yields eight possible representations [11]. However, the specific phenomenology of the $D^{0}\to K_{S}^{0}K_{S}^{0}$ decay imposes stringent theoretical and experimental constraints, which eliminate the majority of these candidates[12, 5].
In order to generate a sizeable NP contribution to the CP asymmetry in this channel, the scalar diquark model must satisfy the following four simultaneous conditions:
  1. 1.
    Simultaneous Coupling to Up- and Down-Type Quarks: The scalar must couple to both up-type and down-type quarks in order to mediate the required $c\to us\bar{s}$ transition through diquark exchange [7].
  2. 2.
    Chirality Mixing: To generate scalar and tensor four-quark operators (in particular $C_{7}$ and $C_{8}$) and enable the chiral interference mechanism responsible for CP violation, the diquark must couple to both left-handed quark doublets $Q_{L}$ and right-handed singlets $u_{R},d_{R}$[7].
  3. 3.
    Flavor Structure: The couplings in the up-quark sector must be symmetric in flavor space to avoid dangerous flavor-changing neutral currents (FCNCs) and to allow the required interference pattern [11].
  4. 4.
    Consistency with Flavor Constraints: The model must respect current experimental bounds, in particular those from neutral meson mixing. Crucially, the diquark must not induce tree-level contributions to $B_{s}-\bar{B}_{s}$ mixing, which severely constrains many diquark and leptoquark scenarios [13].
Imposing these conditions directs us to scalar fields transforming as weak isospin singlets with hypercharge $Y=1/3$, which can exist in two viable color representations: the color anti-triplet and the color sextet [11]. These correspond to scalar fields transforming as:
$\phi_{3}\sim(\mathbf{\bar{3}},\mathbf{1},1/3)_{SM},\quad\text{and}\quad\phi_{6}\sim(\mathbf{6},\mathbf{1},1/3)_{SM}$ (10)
under $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$.Unlike other representations which couple exclusively to right-handed quarks (e.g. $d_{R}d_{R}$) or purely to left-handed doublets, these representations allow simultaneous couplings to both chiralities. The corresponding interaction Lagrangian can be generalized for both cases as:
$\mathcal{L}=\left(\frac{1}{2}\tilde{\lambda}_{ij}^{L}(\overline{Q}_{i}^{I,\alpha})^{c}Q_{j}^{J,\beta}\epsilon_{IJ}+\tilde{\lambda}_{ij}^{R}\overline{u}_{i}^{\alpha c}d_{j}^{\beta}\right)\Gamma_{\alpha\beta}^{n}\phi_{n}+\text{h.c.},$ (11)
where $\alpha,\beta,n$ are color indices, $I,J$ are $SU(2)_{L}$ indices, and $i,j$ are flavor indices. The tensor $\Gamma_{\alpha\beta}^{n}$ dictates the color topology: for the anti-triplet $\phi_{3}$, it is proportional to the totally antisymmetric tensor $\epsilon_{\alpha\beta n}$, whereas for the sextet $\phi_{6}$, it represents a symmetric color tensor.The simultaneous presence of $\tilde{\lambda}^{L}$ and $\tilde{\lambda}^{R}$ is the key feature enabling the chiral interference necessary for CP violation, which is absent in the SM. The symmetric color structure of the sextet representation avoids the destructive color suppression inherent to the antisymmetric triplet.
IV Phenomenology of Scalar Diquarks in $D^{0}\to K_{S}^{0}K_{S}^{0}$
In the SM, the decay $D^{0}\to K_{S}^{0}K_{S}^{0}$ is strongly suppressed compared to $D^{0}\to K^{+}K^{-}$. In the exact U-spin symmetry limit, corresponding to the interchange $s\leftrightarrow d$, the amplitude for $D^{0}\to K^{0}\bar{K}^{0}$ vanishes due to the cancellation of leading contributions [4, 14]. As a consequence, the SM predicts a characteristic anti-correlation between the direct CP asymmetries of the charged kaon and pion modes,
$a_{CP}^{\rm dir}(D^{0}\to K^{+}K^{-})+a_{CP}^{\rm dir}(D^{0}\to\pi^{+}\pi^{-})\approx 0,$ (12)
up to corrections induced by U-spin breaking [4, 15].
The most stringent test of Eq. (12) is provided by the LHCb determination of the direct CP asymmetries extracted from a global combination of time-integrated measurements, including $\Delta A_{CP}$, using Run 1 and Run 2 data [6]. The reported values are
$a_{CP}^{\rm dir}(D^{0}\to K^{+}K^{-})$ $=(7.7\pm 5.7)\times 10^{-4},$ (13)
$a_{CP}^{\rm dir}(D^{0}\to\pi^{+}\pi^{-})$ $=(23.2\pm 6.1)\times 10^{-4}.$ (14)
Their sum is
$a_{CP}^{\rm dir}(K^{+}K^{-})+a_{CP}^{\rm dir}(\pi^{+}\pi^{-})=(30.8\pm 11.4)\times 10^{-4},$ (15)
which deviates from zero at the level of approximately $2.7\sigma$ [6]. Notably, both asymmetries are positive, in contrast with the leading-order U-spin prediction.
A more general U-spin relation involving neutral kaons can be obtained by decomposing the final states into U-spin singlet and triplet components [4, 16],
$A(D^{0}\to K^{+}K^{-})+A(D^{0}\to\pi^{+}\pi^{-})+\sqrt{2}\,A(D^{0}\to K_{S}^{0}K_{S}^{0})=0,$ (16)
which highlights the special role of the $D^{0}\to K_{S}^{0}K_{S}^{0}$ channel in probing U-spin breaking and potential new physics effects [14].
In the SM, the decay $D^{0}\to K_{S}^{0}K_{S}^{0}$ proceeds dominantly through W-exchange topologies, whose leading contribution vanishes in the exact U-spin limit. The amplitude therefore arises only at first order in U-spin breaking,
$A_{\rm SM}(D^{0}\to K_{S}^{0}K_{S}^{0})\;\sim\;\epsilon_{U}\,A_{SU(3)},\qquad\epsilon_{U}\sim\frac{m_{s}-m_{d}}{\Lambda_{\rm QCD}},$ (17)
resulting in a highly suppressed branching ratio and CP asymmetry [4, 5]. This suppression makes the mode particularly sensitive to new physics contributions.
Experimentally, LHCb has measured
$A_{CP}(D^{0}\to K_{S}^{0}K_{S}^{0})=(1.86\pm 0.23\pm 0.11)\%,$ (18)
which constitutes the largest CP asymmetry observed in the charm sector and is difficult to accommodate within the SM [LHCb:2023ksks]. New physics contributions transforming purely as $\Delta U=0$ would preserve the anti-correlation
$A_{CP}(D^{0}\to K^{+}K^{-})=-A_{CP}(D^{0}\to\pi^{+}\pi^{-}),$
and are therefore disfavored by Eq. (15). In contrast, operators transforming as $|\Delta U|=1$ naturally violate the U-spin sum rule and can generate the observed pattern [4].
A well-motivated realization of such operators arises from scalar diquarks transforming as either a color sextet ($S_{6}$) or color triplet ($S_{3}$), whose flavor-dependent couplings generate effective four-quark interactions [7, 17],
$C_{1}^{\rm NP}(q)\;\simeq\;\frac{\lambda_{uq}\lambda_{cq}^{*}}{M_{\phi}^{2}},\qquad q=d,s.$ (19)
A hierarchy $\lambda_{ud}>\lambda_{us}$ naturally enhances CP violation in $D^{0}\to K_{S}^{0}K_{S}^{0}$ while simultaneously reproducing the observed pattern in charged final states.
The effective Hamiltonian governing the transitions $c\bar{u}\to\bar{q}q$ ($q=d,s$) can be written as [18, 7]
$\mathcal{L}_{\text{eff}}=-\sum_{q=d,s}\sum_{X}N_{SM}\left(C_{X}^{q}Q_{X}^{q}+C_{X}^{\prime q}Q_{X}^{\prime q}\right),$ (20)
where
$N_{SM}=\frac{4G_{F}}{\sqrt{2}}V_{uq}V_{cq}^{*},$ (21)
The complete set of dimension-six four-quark operators relevant for $c\to uq\bar{q}$ transitions is
$Q_{1}^{q}$ $=(\bar{u}^{\alpha}\gamma_{\mu}P_{L}q^{\beta})(\bar{q}^{\beta}\gamma^{\mu}P_{L}c^{\alpha}),$ (22)
$Q_{2}^{q}$ $=(\bar{u}^{\alpha}\gamma_{\mu}P_{L}q^{\alpha})(\bar{q}^{\beta}\gamma^{\mu}P_{L}c^{\beta}),$
$Q_{3}^{q}$ $=(\bar{u}^{\alpha}P_{L}q^{\beta})(\bar{q}^{\beta}P_{L}c^{\alpha}),$
$Q_{4}^{q}$ $=(\bar{u}^{\alpha}P_{L}q^{\alpha})(\bar{q}^{\beta}P_{L}c^{\beta}),$
$Q_{5}^{q}$ $=(\bar{u}^{\alpha}\sigma_{\mu\nu}P_{L}q^{\beta})(\bar{q}^{\beta}\sigma^{\mu\nu}P_{L}c^{\alpha}),$
$Q_{6}^{q}$ $=(\bar{u}^{\alpha}\sigma_{\mu\nu}P_{L}q^{\alpha})(\bar{q}^{\beta}\sigma^{\mu\nu}P_{L}c^{\beta}),$
$Q_{7}^{q}$ $=(\bar{u}^{\alpha}P_{R}q^{\beta})(\bar{q}^{\beta}P_{R}c^{\alpha}),$
$Q_{8}^{q}$ $=(\bar{u}^{\alpha}P_{R}q^{\alpha})(\bar{q}^{\beta}P_{R}c^{\beta}),$
$Q_{9}^{q}$ $=(\bar{u}^{\alpha}\sigma_{\mu\nu}P_{R}q^{\beta})(\bar{q}^{\beta}\sigma^{\mu\nu}P_{R}c^{\alpha}),$
$Q_{10}^{q}$ $=(\bar{u}^{\alpha}\sigma_{\mu\nu}P_{R}q^{\alpha})(\bar{q}^{\beta}\sigma^{\mu\nu}P_{R}c^{\beta}),$
with primed operators obtained by $P_{L}\leftrightarrow P_{R}$. These operators can be classified according to their Lorentz and color structure:
  • Current–current operators:
    $Q_{1}^{q},\quad Q_{2}^{q},$
    which dominate in the SM.
  • Scalar operators:
    $Q_{3}^{q},\quad Q_{4}^{q},\quad Q_{7}^{q},\quad Q_{8}^{q}.$
  • Tensor operators:
    $Q_{5}^{q},\quad Q_{6}^{q},\quad Q_{9}^{q},\quad Q_{10}^{q}.$
In the SM, only the operators $Q_{1,2}^{q}$ receive sizable contributions at the electroweak scale. The scalar and tensor operators are either absent or highly suppressed, resulting in small CP asymmetries in charm decays.
In contrast, scalar diquark exchange generates new contributions at tree level. Integrating out the scalar diquark $\phi$ yields
$C_{1}^{q}$ $=\pm C_{2}^{q}=-\frac{\lambda_{uq}^{L*}\lambda_{cq}^{L}}{2N_{q}N_{SM}M_{\phi}^{2}},$ (23)
$C_{7}^{q}$ $=-C_{8}^{q}=-4C_{9}^{q}=4C_{10}^{q}=\frac{\lambda_{uq}^{L*}\lambda_{cq}^{R}}{2N_{q}N_{SM}M_{\phi}^{2}},$ (24)
while the remaining operators are not generated independently at tree level.
The sign and magnitude of the Wilson coefficients depend crucially on the color representation of the diquark:
  • Color sextet ($\mathbf{\bar{6}}$):
    $C_{1}^{NP}=C_{2}^{NP}.$
    This symmetric color structure allows constructive interference between color contractions, enhancing exchange amplitudes such as $D^{0}\to K_{S}^{0}K_{S}^{0}$.
  • Color triplet ($\mathbf{3}$):
    $C_{1}^{NP}=-C_{2}^{NP}.$
    The antisymmetric color structure leads to destructive interference, significantly suppressing its contribution.
This distinction has important phenomenological consequences. The color–sextet diquark efficiently generates $\Delta U=0$ and $\Delta U=1$ operators and can substantially enhance CP violation in exchange–dominated channels. In contrast, the color–triplet diquark primarily contributes to $\Delta U=1$ operators and produces smaller effects due to color suppression.
The presence of new scalar and tensor operators, together with the enhanced exchange amplitude in the sextet case, provides a natural mechanism for generating large CP asymmetries in $D^{0}\to K_{S}^{0}K_{S}^{0}$ while simultaneously affecting other charm decay observables.
IV.1 Scalar diquark contribution to direct CP violation in $D^{0}\to K_{S}^{0}K_{S}^{0}$
IV.1.1 Color-sextet case
In the color-sextet scalar diquark scenario, the relation
$C_{1}^{\rm NP}=C_{2}^{\rm NP}$ (25)
implies that the dominant new physics contributions arise from the current–current operators $O_{1,2}$. Integrating out a scalar diquark of mass $M_{\Phi}$ generates effective four-quark operators contributing to the transitions $c\to u\bar{q}q$ ($q=d,s$),
$\mathcal{H}_{\rm eff}^{\rm NP}=\frac{\lambda_{uq}\lambda_{cq}^{*}}{16M_{\Phi}^{2}}\left(O_{1}^{q}+O_{2}^{q}\right).$ (26)
The total decay amplitude for $D^{0}\to K_{S}^{0}K_{S}^{0}$ receives both SM and new physics contributions and can be written as
$\mathcal{A}_{\rm total}=\lambda_{sd}A_{sd}^{\rm SM}+\lambda_{b}A_{b}^{\rm SM}+\mathcal{A}_{H},$ (27)
where
$\lambda_{sd}=V_{cs}V_{us}^{*},\qquad\lambda_{b}=V_{cb}V_{ub}^{*},$ (28)
and $\mathcal{A}_{H}$ denotes the diquark-induced contribution.
For a scalar diquark of mass $M_{\Phi}$, the new physics amplitude is
$\mathcal{A}_{H}=\frac{1}{8M_{\Phi}^{2}}\left[\lambda_{us}^{L}\lambda_{cs}^{L*}\langle K_{S}^{0}K_{S}^{0}|O_{H}^{s}|D^{0}\rangle-\lambda_{ud}^{L}\lambda_{cd}^{L*}\langle K_{S}^{0}K_{S}^{0}|O_{H}^{d}|D^{0}\rangle\right].$ (29)
The direct CP asymmetry is defined as
$A_{CP}^{\rm dir}=\frac{|\mathcal{A}|^{2}-|\bar{\mathcal{A}}|^{2}}{|\mathcal{A}|^{2}+|\bar{\mathcal{A}}|^{2}},$ (30)
which can be expressed as
$A_{CP}^{\rm dir}=\frac{2\left|\frac{\lambda_{b}}{\lambda_{sd}}\right|{\rm Im}\left(\frac{A_{b}}{A_{sd}}\right)\sin\gamma+{\rm Im}(\Delta_{\rm NP})}{1+|R_{sd}|^{2}},$ (31)
where
  • $R_{sd}\equiv\mathcal{A}_{H}/(\lambda_{sd}A_{sd}^{\rm SM})$ measures the relative size of the new physics contribution,
  • $\gamma$ is the CKM weak phase,
  • ${\rm Im}(A_{b}/A_{sd})$ encodes the SM strong phase difference,
  • the new physics contribution is given by
    ${\rm Im}(\Delta_{\rm NP})=2r\sin(\phi_{\rm NP})\sin(\delta),$ (32)
    with
    $r=\left|\frac{\mathcal{A}_{H}}{\lambda_{sd}A_{sd}^{\rm SM}}\right|,$ (33)
    $\phi_{\rm NP}$ the new weak phase from the diquark couplings, and $\delta$ the relative strong phase between SM and new physics amplitudes.
Since rescattering effects are expected to be significant in $D^{0}\to K_{S}^{0}K_{S}^{0}$, large strong phases are plausible [3, 8, 19]. Assuming $r\sim 0.01$ and maximal weak and strong phases, the resulting CP asymmetry can reach
$A_{CP}^{\rm dir}\sim 2r\sim 2\%,$ (34)
consistent with current experimental measurements.
A summary of the predicted CP asymmetry as a function of the diquark mass is shown in Fig. 1.
Figure 1: Direct CP asymmetry $a_{CP}^{dir}$ in $D^{0}\to K_{S}^{0}K_{S}^{0}$ vs. diquark mass $M_{\Phi}$. The predictions for the color-sextet ($S_{6}$, solid blue) and color-triplet ($S_{3}$, dashed green) models are compared against the LHCb measurement. A suppression factor of $0.1$ is applied to the $S_{3}$ curve to account for the destructive interference arising from the relation $C_{1}^{q}=-C_{2}^{q}$.
IV.1.2 Color-triplet case
In the color-triplet diquark scenario, the antisymmetric color structure implies
$C_{1}^{\rm NP}=-C_{2}^{\rm NP},$ (35)
which suppresses the contribution from current–current operators due to destructive interference. In addition, scalar and tensor operators ($O_{7-10}$) generated by the diquark provide an additional source of CP violation.
The resulting direct CP asymmetry can be written as
$A_{CP}^{\rm dir}=r_{T}\sin(\phi_{\rm NP}^{LL})\sin(\delta_{T})+r_{P}\sin(\phi_{\rm NP}^{LR})\sin(\delta_{P}),$ (36)
where the two terms correspond to contributions from left-handed and chirality-flipping operators, respectively.
The relative amplitudes are given by
$r_{T}$ $=\left|\frac{\mathcal{A}_{H}}{\lambda_{sd}A_{sd}^{\rm SM}}\right|=\left|\frac{\lambda_{us}^{L}\lambda_{cs}^{L*}}{4M_{\Phi}^{2}}\right|\frac{1}{\frac{G_{F}}{\sqrt{2}}|V_{cs}V_{us}^{*}|}\left|\frac{\langle K_{S}^{0}K_{S}^{0}|O_{2}-O_{1}|D^{0}\rangle}{\langle K_{S}^{0}K_{S}^{0}|\mathcal{H}_{\rm SM}|D^{0}\rangle}\right|,$ (37)
$r_{P}$ $\approx\frac{1}{\frac{G_{F}}{\sqrt{2}}|V_{cs}V_{us}^{*}|}\left|\frac{\lambda_{uq}^{L*}\lambda_{cq}^{R}}{4M_{\Phi}^{2}}\right|\frac{\langle K_{S}^{0}K_{S}^{0}|Q_{EWP}|D^{0}\rangle}{\langle K_{S}^{0}K_{S}^{0}|\mathcal{H}_{\rm SM}|D^{0}\rangle}.$ (38)
The phases
$\phi_{\rm NP}^{LL}={\rm Arg}(\lambda_{us}^{L}\lambda_{cs}^{L*}),\qquad\phi_{\rm NP}^{LR}={\rm Arg}(\lambda_{us}^{L}\lambda_{cs}^{R*}),$ (39)
represent the new weak phases, while $\delta_{T,P}$ denote strong phases generated by final-state interactions.
Assuming perturbative couplings $\lambda_{ij}^{L,R}\sim 0.1$ and a diquark mass of order 1 TeV, both contributions are parametrically similar. However, due to destructive color interference in the triplet case, the overall CP asymmetry remains significantly smaller than in the sextet scenario.
A summary of the predicted CP asymmetry as a function of the diquark mass is shown in Fig. 1.
IV.1.3 Compatibility with $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$
An important consistency check of the color-sextet diquark scenario comes from its impact on the singly Cabibbo-suppressed modes $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$. For a representative new physics amplitude ratio $r\sim 1\%$, one can estimate the induced CP asymmetry using the framework of Ref. [7]. In particular, for the $K^{+}K^{-}$ channel, one obtains
$A_{CP}^{\rm dir}(D^{0}\to K^{+}K^{-})\approx 0.8\%,$ (40)
which is significantly larger than the SM expectation and within the range suggested by current experimental measurements.
Since the diquark couplings are, in general, flavor-dependent, they naturally generate $|\Delta U|=1$ operators, thereby violating the U-spin sum rule discussed previously. In particular, a hierarchy
$\lambda_{ud}>\lambda_{us},$ (41)
can simultaneously reproduce the observed positive pattern of CP asymmetries in both $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ channels, while also enhancing the asymmetry in $D^{0}\to K_{S}^{0}K_{S}^{0}$.
The sensitivity of different decay channels depends strongly on their underlying topologies. The decay $D^{0}\to K_{S}^{0}K_{S}^{0}$ is dominated by exchange ($E$) topologies and is exceptionally sensitive to new physics, since the leading tree-level amplitude vanishes in the U-spin limit. In contrast, the decays $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$ are dominated by color-allowed $W$-emission ($T^{\prime}$) topologies.
In these emission-dominated channels, the color-sextet diquark, which satisfies $C_{1}^{\rm NP}=C_{2}^{\rm NP}$, enhances the emission amplitude through constructive interference without color suppression. This leads to sizable CP asymmetries while preserving consistency with other charm observables.
Figure 2: A comprehensive multi-channel comparison showing the absolute asymmetries for $D^{0}\to K_{S}^{0}K_{S}^{0}$, $D^{0}\to K^{+}K^{-}$, and $D^{0}\to\pi^{+}\pi^{-}$ under both the $S_{6}$ (left panel) and $S_{3}$ (right panel) scenarios.
Figure 2 illustrates the characteristic hierarchy of CP asymmetries induced by the sextet diquark across different decay modes:
  • $D^{0}\to K_{S}^{0}K_{S}^{0}$: This channel exhibits the strongest sensitivity to new physics. Because the leading SM amplitude is suppressed in the U-spin limit, the diquark contribution can dominate, leading to asymmetries in the range $0.5\%$$1.5\%$ for a diquark mass around 1 TeV.
  • $D^{0}\to K^{+}K^{-}$: This channel is dominated by color-allowed emission. The diquark interferes with the SM amplitude, producing asymmetries of order $0.7\%$$0.8\%$.
  • $D^{0}\to\pi^{+}\pi^{-}$: This decay exhibits similar behavior, with asymmetries of comparable magnitude. The flavor hierarchy $\lambda_{ud}>\lambda_{us}$ naturally allows both kaon and pion asymmetries to be positive, in agreement with the observed violation of the U-spin sum rule.
Table 2: Comparison of predicted direct CP asymmetries in the $S_{6}$ diquark model (for $M_{S_{6}}\sim 1$ TeV) with Standard Model expectations and experimental measurements.
Decay Channel SM Expectation Sextet Model Experimental Result
$D^{0}\to K_{S}^{0}K_{S}^{0}$ $\lesssim 10^{-4}$ $0.5\%$$1.5\%$ $(1.86\pm 1.12)\%$
$D^{0}\to K^{+}K^{-}$ $\lesssim 10^{-4}$ $0.7\%$$0.8\%$ $(7.7\pm 5.7)\times 10^{-4}$
$D^{0}\to\pi^{+}\pi^{-}$ $\lesssim 10^{-4}$ $0.6\%$$0.7\%$ $(23.2\pm 6.1)\times 10^{-4}$
$\Delta A_{CP}$ $\approx 0$ $0.1\%$ to $0.6\%$ $(0.645\pm 0.180)\%$
V Conclusion
We have investigated the phenomenological effects of scalar diquarks on CP violation in the decay channels $D^{0}\to K_{S}^{0}K_{S}^{0},K^{+}K^{-},\pi^{+}\pi^{-}$. It is well-known that the expected Standard Model asymmetries in these channels remain below the percent level. We have shown that the introduction of a $S_{6}$ scalar diquark with a mass of order 1 TeV and perturbative couplings can naturally generate CP asymmetries at the percent level, consistent with current experimental observations for $D^{0}\to K_{S}^{0}K_{S}^{0}$.
A comparative analysis of $S_{6}$ and $S_{3}$ scalar diquark representations demonstrates that the $S_{6}$ scalar diquark provides the most viable explanation of the data. Its symmetric color structure, characterized by the relation $C_{1}^{\rm NP}=C_{2}^{\rm NP}$, avoids the color suppression present in exchange amplitudes with $S_{3}$ scalar diquark. As a consequence, it allows for constructive interference with the SM contribution. This leads to an enhancement of the $D^{0}\to K_{S}^{0}K_{S}^{0}$ decay amplitude and of the corresponding CP asymmetry. In contrast, the color-triplet representation, which implies $C_{1}^{\rm NP}=-C_{2}^{\rm NP}$, induce destructive color interference and its effect on $D^{0}\to K_{S}^{0}K_{S}^{0}$ CP asymmetry is suppressed.
Moreover, the flavor-dependent structure of the diquark couplings, which allows to have the hierarchy $\lambda_{ud}>\lambda_{us}$, provides a coherent description of CP violation across multiple charm decay channels. This hierarchy enhances the asymmetry in $D^{0}\to K_{S}^{0}K_{S}^{0}$ while reproducing the observed pattern of positive direct CP asymmetries in $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$. In addition, it can generate the observed deviation from the U-spin sum rule.
As a conclusion, the $S_{6}$ scalar diquarks are a well-motivated candidates for explaining enhanced CP violation in charm decays. Such states arise naturally in grand unified theories based on gauge groups such as $SO(10)$ and $E_{6}$. Future precision measurements of CP asymmetries in charm decays, as well as direct searches for scalar diquarks at high-energy colliders, will provide important tests of this scenario and further clarify the origin of CP violation in the charm sector.
Acknowledgements
We dedicate this work to the memory of our esteemed colleague, Dr. Gaber Faisel, who sadly passed away at the beginning of this project. We are deeply grateful for his foundational insight regarding the diquark contributions to the $D\to K_{S}K_{S}$ decay, which motivated this research. The work of S. K. is partially supported by Science, Technology &\& Innovation Funding Authority (STDF) under grant number 48173. The work of D.D. is supported by Secretaria de Ciencia, Humanidades, Tecnologia e Innovación (SECIHTI) and Sistema Nacional de Investigadoras e Investigadores (S.N.I.I.), Mexico.